How do I solve the system in matrices?:
{x+y=2
{x-y=-4
{2x+y=1
and
{x+y=4
{2x-y=2

Answer by Alan3354(69443)

How do I solve the system in matrices?:
{x+y=2
{x-y=-4
{2x+y=1
You have 3 eqns in x & y ??
{x+y=2
{x-y=-4
---------
2x = -2
x = -1
-------
y = 3
---------
You can solve an augmented 2 by 2 matrix with any 2 of the 3 equations.
If you assume 0z for each and make it 3 by 3, z is undefined anyway.
------------------
and
{x+y=4
{2x-y=2
----------
|1 1 -4|
|2-1 -2|
Det = -1-2 = -3
---------
Dx = -2 -4 = -6
x = Dx/Det = 2
----------
-Dy = -2 + 8 = 6
y = -6/Det = 2
----------------
x = 2
y = 2